# On the Limit of Language Models as Planning Formalizers

### v1

@@ Appendix B Prompts
For the LLM-as-planner, we give all the models the following prompt for BlocksWorld-111:
@@
Appendix C Results for Templated BlocksWorld-111
@@
gemma-2-9b-it 0/111 0/111
gemma-2-9$\dagger$ - 5/111
gemma-2-27b-it 21/111 12/111
gemma-2-27$\dagger$ - 3/111
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gpt-3.5-turbo 14/111 4/111
gpt-4o-mini 9/111 5/111
gpt-4$\dagger$ - 7/111
gpt-4o 87/111 76/111
gpt-4$\dagger$ - 37/111
@@
gemma-2-9b-it 71/111 10/111
gemma-2-9$\dagger$ - 5/111
gemma-2-27b-it 90/111 89/111
gemma-2-27$\dagger$ - 9/111
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gpt-3.5-turbo 41/111 33/111
gpt-4o-mini 71/111 64/111
gpt-4$\dagger$ - 1/111
gpt-4o 99/111 99/111
gpt-4$\dagger$ - 32/111
Table 5: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on Moderately Templated and Heavily Templated BlocksWorld-111.
Table 5 displays results for Heavily Templated and Moderately Templated BlocksWorld-111.
Appendix D Results for Heavily Templated MysteryBlocksWorld-100
@@
gpt-4$\dagger$ - 0/100
Table 6: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Heavily Templated MysteryBlocksWorld-100.
Appendix E Sample Model Output
@@
Appendix F Related Works Comparison

### v2

@@ Appendix B Prompts
For the LLM-as-planner, we give all the models the following prompt for BlocksWorld-100:
@@
Appendix C Detailed Results
Beyond the visualizations above, we show the detailed results of all models on all simulations of all naturalness levels.
C.1 Results for BlocksWorld-100
@@
Natural
Models Solvability Correctness
gemma-2-9b-it 3/100 3/100
gemma-2-9b-itp - 9/100
gemma-2-27b-it 0/100 0/100
gemma-2-27b-itp - 11/100
Llama-3.1-8B 0/100 0/100
Llama-3.1-8Bp - 1/100
Llama-3.1-70B 0/100 0/100
Llama-3.1-70Bp - 13/100
gpt-3.5-turbo 2/100 1/100
gpt-4o-mini 19/100 3/100
gpt-4o-minip - 7/100
gpt-4o 64/100 60/100
gpt-4op - 33/100
o1-preview 91/100 91/100
o1-previewp - 82/100
o3-mini 79/100 68/100
o3-minip - 87/100
@@
gemma-2-9b-it 0/100 0/100
gemma-2-9b-itp - 5/100
gemma-2-27b-it 17/100 10/100
gemma-2-27b-itp - 3/100
Llama-3.1-8B 0/100 0/100
Llama-3.1-8Bp - 1/100
Llama-3.1-70B 0/100 0/100
Llama-3.1-70Bp - 10/100
gpt-3.5-turbo 14/100 4/100
gpt-4o-mini 9/100 5/100
gpt-4o-minip - 7/100
gpt-4o 77/100 67/100
gpt-4op - 35/100
o3-mini 82/100 70/100
o3-minip - 87/100
@@
gemma-2-9b-it 61/100 10/100
gemma-2-9b-itp - 5/100
gemma-2-27b-it 81/100 80/100
gemma-2-27b-itp - 7/100
Llama-3.1-8B 0/100 0/100
Llama-3.1-8Bp - 0/100
Llama-3.1-70B 0/100 0/100
Llama-3.1-70Bp - 10/100
gpt-3.5-turbo 39/100 29/100
gpt-4o-mini 66/100 59/100
gpt-4o-minip - 1/100
gpt-4o 89/100 89/100
gpt-4op - 29/100
o3-mini 94/100 94/100
o3-minip - 96/100
Table 3: Performance of LLM-as-formalizer and LLM-as-planner (p) all BlocksWorld-100 data.
Table 3 displays results for all results for BlocksWorld-100.
C.2 Results for MysteryBlocksWorld-100
@@
gpt-4op - 0/100
o3-mini 95/100 95/100
o3-minip - 74/100
Table 4: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated MysteryBlocksWorld-100.
C.3 Results for Logistics-100
Table 5 displays results for all naturalness settings for Logistics-100
Metrics
Natural
Models Solvability Correctness
gemma-2-9b-it 1/100 0/100
gemma-2-9b-itp - 0/100
gemma-2-27b-it 1/100 0/100
gemma-2-27b-itp - 0/100
Llama-3.1-8B 0/100 0/100
Llama-3.1-8Bp - 1/100
Llama-3.1-70B 0/100 0/100
Llama-3.1-70Bp - 0/100
gpt-3.5-turbo 1/100 0/100
gpt-3.5-turbop - 0/100
gpt-4o-mini 13/100 0/100
gpt-4o-minip - 0/100
gpt-4o 20/100 2/100
gpt-4op - 1/100
o3-mini 43/100 7/100
o3-minip - 14/100
Moderately Templated
Models Solvability Correctness
gemma-2-9b-it 0/100 0/100
gemma-2-9b-itp - 0/100
gemma-2-27b-it 3/100 0/100
gemma-2-27b-itp - 0/100
Llama-3.1-8B 0/100 0/100
Llama-3.1-8Bp - 0/100
Llama-3.1-70B 0/100 0/100
Llama-3.1-70Bp - 0/100
gpt-3.5-turbo 0/100 0/100
gpt-3.5-turbop - 0/100
gpt-4o-mini 29/100 0/100
gpt-4o-minip - 0/100
gpt-4o 34/100 13/100
gpt-4op - 0/100
o3-mini 51/100 39/100
o3-minip - 21/100
Heavily Templated
Models Solvability Correctness
gemma-2-9b-it 1/100 0/100
gemma-2-9b-itp - 0/100
gemma-2-27b-it 0/100 0/100
gemma-2-27b-itp - 1/100
Llama-3.1-8B 0/100 0/100
Llama-3.1-8Bp - 0/100
Llama-3.1-70B 0/100 0/100
Llama-3.1-70Bp - 0/100
gpt-3.5-turbo 1/100 1/100
gpt-3.5-turbop - 0/100
gpt-4o-mini 26/100 0/100
gpt-4o-minip - 0/100
gpt-4o 33/100 13/100
gpt-4op - 0/100
o3-mini 55/100 47/100
o3-minip - 21/100
Table 5: Performance of LLM-as-formalizer and LLM-as-planner (p) all Logistics-100 data.
C.4 Results for Barman-100
Figure 5: Performance for Barman-100.
Table 6 displays results for Barman-100.
Models Solvability Correctness
Llama-3.1-8B 0/100 0/100
Llama-3.1-8Bp - 0/100
Llama-3.1-70B 0/100 0/100
Llama-3.1-70Bp - 15/100
gemma-2-9b-it 0/100 0/100
gemma-2-9b-itp - 2/100
gemma-2-27b-it 0/100 0/100
gemma-2-27b-itp - 0/100
gpt-3.5-turbo 0/100 0/100
gpt-3.5-turbop - 4/100
gpt-4o-mini 0/100 0/100
gpt-4o-minip - 8/100
gpt-4o 1/100 0/100
gpt-4op - 3/100
o3-mini 1/100 0/100
o3-minip - 19/100
Table 6: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated Barman-100.
Appendix D Sample Model Output
@@
Appendix E Related Works Comparison
@@
Appendix F Methodology Comparison
Both methodologies used in this paper incorporate LLMs into planning. There are pros and cons to each methodology. When using LLM-as-planner, we have a lightweight solution that returns results very quickly. However, due to the lack of reasoning skills in LLMs, they often struggle to create formal plans. Meanwhile using LLM-as-formalizer provides better executability and interpretability, though it uses a solver, which may results in getting results slower. We believe that for performance reasons that LLM-as-formalizer is the superior methodology.